OpenStudy (anonymous):

Could you please help with this problem write as a single fraction -2+7a+2b/6-a+4b/8

OpenStudy (anonymous):

can u either draw it or put parentheses

OpenStudy (anonymous):

(-2+7a+2b/6)-a+4b/8 it is a fraction

OpenStudy (anonymous):

-2+7a+2b/6-a+4b/8

OpenStudy (anonymous):

From the problem expression:\[7 a-a+\frac{2 b}{6}+\frac{4 b}{8}-2=6 a+\frac{5 b}{6}-2 \]

OpenStudy (anonymous):

http://en.cppreference.com/w/cpp/language/operator_precedence Refer to lines 5 and 6.

OpenStudy (anonymous):

multiply and divide have stronger binding power then add and subtract.